OpenStudy (anonymous):

(-3y2 - 7y - 9) - (4y2 + 6y + 9) A. y2 - y B. -9y2 - 11y - 18 C. 7y2 + 13y + 18 D. -7y2 - 13y - 18

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

D is the answer, just combine the like terms

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Evaluate the polynomial 6x - y for x = - 3 and y = 2. A. 15 B. –20 C. 4 D. –16

OpenStudy (anonymous):

substitute the x by -3 and y by 2 6*-3 - 2=?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

You have checked out b books from the library, where b is a number greater than 5. You plan to read three of the books and return them at the end of this week. In how many ways can you pick the books you will read first, second, and third? A. b3 ways B. (b3 - 3b2 + 2b) ways C. (b3 - 9b2 + 27b - 27) ways D. (5b - 3) ways

OpenStudy (anonymous):

-18-2= -20

OpenStudy (anonymous):

cmon cupcape, try working these out yourself :p